### v1

@@ 2.1 Plane fitting
Once the test satellites are distributed around their host galaxy, the next step involves performing plane fitting to analyze the spatial arrangement of these satellites. This technique helps us understand whether the satellites are arranged in a planar structure or scattered more isotropically around the host galaxy. For this purpose, a viable method of unweighted plane-fitting technique is used Metz et al. (2007). This method commences with the calculation of the moment of inertia tensor of the satellite distribution, followed by its diagonalization. The process begins by determining the centroid, $r_{0}$, of the data points:
$r_{0}=\frac{1}{N}\sum_{i=1}^{N}r_{i}$ (5)
Subsequent to this, an eigenvalue analysis of the moment of inertia tensor, $\mathbf{T}_{0}$, is conducted for the position vectors, $\hat{r}_{i}=r_{i}-r_{0},i=1\ldots n$ (Metz et al., 2007; Pawlowski et al., 2013).
$T_{0}=\sum_{i=1}^{N}\left[(\mathbf{r}_{i}-\mathbf{r}_{0})^{2}\cdot\mathbf{1}-(\mathbf{r}_{i}-\mathbf{r}_{0})\cdot(\mathbf{r}_{i}-\mathbf{r}_{0})^{T}\right]$ (6)
Here 1 denotes the unit matrix and $\mathbf{r}^{\textbf{T}}$ is the transposed version of the vector $\mathbf{r}$. The square root of the eigenvalues of the moment of inertia indicates the extent along three axes (a, b, c) of the fitted ellipsoid to the satellite distribution. These values are proportional to the rms deviation relative to the eigenvectors of T. The eigenvector corresponding to the largest eigenvalue defines the normal of the plane, encompassing the centroid. Meanwhile, the eigenvectors corresponding to the intermediate and smallest eigenvalues indicate the directions of the intermediate and major axes of the distribution, respectively. Consequently, the axial ratios c/a and b/a are determined. A small value of c/a suggests two possibilities: if b/a is large, it indicates an oblate distribution resembling a thin plane; if b/a is similarly small ($c/a\approx b/a$), it suggests a narrow prolate distribution resembling a filament-like shape (Metz et al., 2007).
Figure 2: Comparison of eccentricity, $(e)$, distributions across different $\theta$: The figure presents the eccentricity CDF for three different $\theta_{max}$: 40°, 60°, and 80°, from left to right, respectively, as shown in gray. The predominant black line in each panel signifies observed data for the Milky Way satellites from (Li et al., 2021), while the overlaid gray CDF represent simulated results for $N_{\text{sat}}$ test satellites.
@@ 2.2 Adopted Potential
Understanding the gravitational potential and mass distribution of the Milky Way is crucial for numerically integrating the orbits of test satellites. Despite extensive research over the past century, the exact distribution of the Milky Way’s mass remains a topic of debate. Current estimates of the total mass of the Milky Way range between $0.5-3\times 10^{12}M_{\odot}$ (Wang et al., 2020), varying based on the methodology and underlying assumptions (Bland-Hawthorn & Gerhard, 2016). Therefore, significant uncertainty surrounds the mass of the Milky Way, necessitating the consideration of these ambiguities in mass distribution.
For this study, we use one fiducial potential for the forward integration, and three distinct gravitational potentials, as detailed in Table 1, to investigate their effects on the evolution and inferred stability of the Plane of Satellite Galaxies in backward integration. This is to evaluate what effect adopting an incorrect Milky Way mass has on the inferred satellite plane’s evolution. To numerically integrate orbits, we rely on Galpy (Bovy, 2015) to investigate the orbital properties of the test satellites. Within the galpy.potential module, there exists a model named MWPotential2014, which represents the Milky Way’s gravitational potential and is obtained by fitting to a wide range of observational data. This is a simplified static potential model, which is composed of the following components:
@@
with a power-law index of $\alpha=1.8$, and a cut-off radius of $r_{c}=1.9$ kpc.
The disk component is modeled using the axisymmetric Miyamoto–Nagai disk (Miyamoto & Nagai, 1975):
@@
The dark matter halo is characterized by the Navarro–Frenk–White (NFW) halo (Navarro, 1996):
@@
MWPotential2014 is scaled such that the circular velocity at a distance of 8 kpc from the galactic center within the disc ($z=0$ kpc) is 220 km/s with a virial mass of about $8\times 10^{11}M_{\odot}$. In this study, we refer to this potential as the Fiducial Potential, $MW_{\text{fiducial}}$. For the remaining two Milky Way potentials, $MW_{\text{fiducial}}$ is adjusted by re-scaling the halo mass to achieve the desired mass distribution. Specifically, the $MW_{\text{low}}$ model is configured with a halo mass 40% less than that of $MW_{\text{fiducial}}$, while $MW_{\text{high}}$ has a halo mass 40% greater than that of $MW_{\text{fiducial}}$, thus providing deeper insights into the influence of mismatching the Milky Way’s mass on the inferred behavior and stability of satellite galaxy planes.
Table 2: Parameters for forward integration models.
@@
111 The table lists the model numbers (FI), the number of satellites ($N_{\text{sat}}$), the maximum angle $\theta_{max}$ in degrees, the gravitational potential model ($\textit{MW}_{fiducial}$), and the number of realizations ($N_{\text{realization}}$) for each forward integration scenario.
@@ 2.3 Forward Integration
We conduct forward integration of the satellite system over a span of 5 Gyr for $N$ = 20 random realizations of the initial setups. We select three distinct maximum tangential velocity angles $\theta_{\text{max}}$ = 40°, 60°, and 80°, for the forward integration as shown in Table 2. It is worth noting that the larger the angle, the more eccentric orbits are possible.
The resulting cumulative distribution functions (CDFs) of eccentricity for the three different $\theta$ values are illustrated in Fig. 2. Furthermore, we overlay as a solid black curve the eccentricity data for the observed Milky Way satellites from (Li et al., 2021) onto the plot, which contains the data of 46 satellite galaxies from Fritz et al. (2018), McConnachie & Venn (2020) and Simon (2019). However, for our comparison, we picked 35 satellite galaxies, as 11 of them had an eccentricity $e\gg 1$. This addition allows us to compare the simulated eccentricities with those of observed Milky Way satellite galaxies. From the comparison, it becomes apparent that $\theta_{\text{max}}$ = 80°closely aligns with the observed eccentricities of satellite galaxies. Each forward integration was conducted utilizing the $MW_{\text{fiducial}}$ potential model. Fig. 2 also shows that the range of CDF for Li et al. (2021) spans approximately 0.2 to 1.0. For the simulated eccentricity of our initial test runs, the range is between 0.0 and 1.0, because circular orbits were also included. To align the simulated eccentricity of orbits with the observed data, values ranging between $\theta$ = -20°and 20°were excluded to ensure that no perfectly circular orbits are generated.
@@ 2.4 Uncertainties
Figure 3: Effect of proper motion uncertainties on the plane of satellite galaxies. Each panel shows $N_{\text{realization}}$ forward integrations as black curves with the mean indicated by a dot-dashed line, and $M_{\text{realization}}$ backward integrations as green curves, also with the mean indicated by a dashed line.
To determine the motion of a satellite galaxy precisely, its position, distance and 3D motion are crucial. Spectroscopy is used to track the motion of a galaxy using the Doppler effect. However, it is essential to note that radial velocity measures only one dimension. To determine how a galaxy moves tangentially across the sky, proper motions are measured. The Hubble Space Telescope (HST) has measured the proper motion of most of the 11 brightest, classical satellite galaxies of the Milky Way (Redd, 2018). Since then, largely due to the advent of Gaia, the availability of proper motion information has significantly improved (Gaia Collaboration et al., 2016b, a). However, highly accurate proper motion measurements remain limited and are available only for the more massive and luminous or nearby satellite galaxies.
According to Gaia DR2, the random proper motion errors for Milky Way satellites range approximately between 0.006 and 0.641 mas yr-1 (Fritz et al., 2018), and the corresponding systematic errors are estimated to be approximately 0.035 to 0.066 mas yr-1 (Gaia Collaboration et al., 2018a, b). In Gaia EDR3, the uncertainties in proper motions, derived from both statistical and systematic errors, have approximately halved(Lindegren et al., 2021; McConnachie & Venn, 2020; Li et al., 2021; Battaglia et al., 2022).
In this study, we investigate four different ranges of proper motion uncertainties: $\epsilon_{\mu}=0.00$ mas yr-1, 0.04 mas yr-1, 0.08 mas yr-1, and 0.12 mas yr-1. We specifically opted for 0.00 mas yr-1 proper motion uncertainties to represent cases where proper motion is measured with the highest accuracy. This condition allows us to examine the evolution of the plane of satellite galaxies under various parameters independent of proper motion errors. We randomly sample these proper motion uncertainties from Gaussian distributions with width $\epsilon_{\mu}$.
For distance uncertainties, we chose two specific values: $\epsilon_{\mathrm{dist}}=0$% and 5%. Distances are crucial factors in understanding the spatial distribution and dynamics of satellite galaxies around the Milky Way, and their effects on the inferred stability of the plane of satellite galaxies. Again, distance errors were sampled from a Gaussian with width $\epsilon_{\mathrm{dist}}$.
@@ 2.5 Backward Integration
After completing the forward integration, the orbital parameters – such as right ascension, declination, proper motion in right ascension and declination, distance, and line-of-sight velocity – are recorded for each test satellite system at the final snapshot of all simulation realizations. They are then mocked-observed by adding uncertainties in proper motions and distance, such that:
@@
$\text{dist}_{\text{new}}$ $=\text{dist}+\epsilon_{\text{dist}}$ (12)
Equation (10) – (12) provides the updated proper motions and distance parameters for each test satellite, incorporating their respective uncertainties. For the values of $\epsilon_{\mu}$, we draw them from a Gaussian distribution. In contrast, for $\epsilon_{\text{dist}}$, a percentage error is drawn from a Gaussian distribution, such that more distant satellites have larger absolute distance uncertainties. Table 4 details the ranges of uncertainties in these parameters. After performing mock observations, the system is integrated backward for 5 Gyr. This process is repeated 30 times for each of the 20 forward integrations, using different randomly generated errors each time, resulting in a total of 600 backward integrations to return the test satellite system to its initial state. This procedure is conducted using all three Milky Way potential models. Each row in Table 4 represents a combination of various parameters that could influence the inferred stability of the satellite plane. Overall, with four proper motion uncertainties, two distance uncertainties, and three potential models, there are 24 model combinations for each $\theta_{max}$. Each model undergoes 600 backward integrations, amounting to a total of 43,200 backward integrations.
@@ 3 Results
In this section we present the main results from the simulation  [^2] GitHub Repository of the code used to generate the results: Plane of Satellite Galaxies. For brevity, and since it is the case most closely resembling the eccentricities of the observed Milky Way satellite galaxies, we only show results for $\theta_{max}=80$. We refer the reader to Appendix A for results for the other $\theta_{\text{max}}$, which are qualitatively similar.
3.1 Effect of Proper Motions Uncertainties
@@
Fig. 3 presents the results focusing on introducing artificial proper motion and distance uncertainties to the test satellite system before backward integrating. These results are divided into four panels, each corresponding to a distinct level of proper motion uncertainty.
@@
In each panel, there is a scatter of curves shown in both gray and green, indicating the $N_{\text{realization}}$ and $M_{\text{realization}}$ numbers for forward and backward integration, respectively. The average value at each time step is shown as a dashed-dotted curve, with arrows pointing to the right and left. In this case, we use $MW_{\text{fiducial}}$ for both forward and backward integration, and we don’t add uncertainties in distance . In the upper left panel, when $\epsilon_{\mu}$ = 0.00 mas yr-1, the mean minor-to-major ratio, $c/a$ obtained from backward integration aligns perfectly well with the mean minor-to-major ratio, ${c/a}$ obtained from the forward integration, as each forward integration curve is over-plotted by the backward integration. This shows a high degree of predictability and confirms that the orbital plane remains stable and unaltered in scenarios where the proper motion is measured with perfect accuracy. It also implies that, at least in the theoretical framework of the simulation, the inherent dynamics of the satellite system are deterministic enough to allow for precise backtracking of its orbital path when initial conditions are precisely known. These results serve as a control scenario against which the other panels, which include some amount of proper motion errors. Model BI-01 from Table 4 shows similar behavior that, even after 3 Gyr of backward integration, the difference between forward and backward, $\Delta_{c/a}=0.0$, and fractional change, $f_{c/a}$, is 1.00, meaning that both forward and backward integration yield the same output when proper motion uncertainties remain at zero.
@@
In the upper right panel, we have Gaia level of systematic uncertainties, $\epsilon_{\mu}$ = 0.04 mas yr-1. Here, a slight uncertainty in proper motion leads to a visible divergence between the mean axial ratios of forward and backward integrations. The mean axial ratio from the backward integration does not align as closely with the forward integration as it did in the first case, when there were no uncertainties. This suggests a decrease in the predictability and inferred stability of the orbital plane with the introduction of proper motion errors. However, the curve here seems to have retained its shape to some degree. The effect of uncertainties is also evident from Model BI-02 in Table 4, which shows that with the introduction of $\epsilon_{\mu}=0.04$ mas yr-1, $\Delta_{c/a}$ shifts from 0.00 to 0.073, and $f_{c/a}$ becomes 1.69. Thus, the inferred past plane width is typically 70% higher than the true one.
@@
Now, increasing the proper motion uncertainties to 0.08 mas yr-1 further diverges the means of backward and forward integration, as shown in the lower left panel. This indicates a further decline in the predictability and stability of the orbital plane as proper motion uncertainty increases. Moreover, Table 4 also validates that in Model BI-03, $\Delta_{c/a}$ shifts yet again from 0.07 to 0.17, and $f_{c/a}$ rises to 2.64, meaning the typical inferred plane width is almost three times larger than the true one.
@@
Furthermore, under extreme proper motion uncertainties, 0.12 mas yr-1, the divergence between the mean forward and mean backward integration, shown by the dashed curves, is the most pronounced. The mean axial ratio from backward integration shows a significant divergence from the forward integration mean, implying a severe reduction in the orbital plane’s predictability and stability with the highest proper motion uncertainty examined.
@@
Moreover, the original shape of the curve is no longer discernible and deviates significantly more from the forward integration curve compared to the other uncertainties examined. This suggests that the system’s actual past evolution cannot be reliably traced under such high uncertainties. It shows that the inferred dynamics of the system are highly sensitive to proper motion uncertainties, making it challenging to draw reliable conclusions about the satellite’s orbital evolution under these conditions. This case is represented by Model BI-04 of Table 4; it shows that $\Delta_{c/a}$ = 0.24 and $f_{c/a}$ = 3.32.

### v2

@@ 2.2 Plane fitting
Once the test satellites were distributed around their host galaxy, the next step was to perform plane fitting to analyze their spatial arrangement. This technique allowed us to determine whether the satellites were arranged in a planar structure or scattered more isotropically around the host galaxy. For this purpose, we used an unweighted plane-fitting technique Metz et al. (2007). This method begins with the calculation of the moment-of-inertia tensor of the satellite distribution, which is followed by its diagonalization. The first step of the process is to determine the centroid $r_{0}$, of the data points:
$r_{0}=\frac{1}{N}\sum_{i=1}^{N}r_{i}.$ (5)
Subsequently, an eigenvalue analysis of the moment of inertia tensor, $\mathbf{T}_{0}$, is performed for the position vectors, $\hat{r}_{i}=r_{i}-r_{0},i=1\ldots n$ (Metz et al., 2007; Pawlowski et al., 2013):
$T_{0}=\sum_{i=1}^{N}\left[(\mathbf{r}_{i}-\mathbf{r}_{0})^{2}\cdot\mathbf{1}-(\mathbf{r}_{i}-\mathbf{r}_{0})\cdot(\mathbf{r}_{i}-\mathbf{r}_{0})^{T}\right].$ (6)
Here, 1 denotes the unit matrix and $\mathbf{r}^{\textbf{T}}$ is the transposed version of the vector $\mathbf{r}$. The square root of the eigenvalues of the moment of inertia indicates the extent along three axes (a, b, c) of the fitted ellipsoid to the satellite distribution. These values are proportional to the rms deviation relative to the eigenvectors of T. The eigenvector corresponding to the largest eigenvalue defines the normal of the plane, encompassing the centroid. Meanwhile, the eigenvectors corresponding to the intermediate and smallest eigenvalues indicate the directions of the intermediate and major axes of the distribution, respectively. Consequently, we determined the axial ratios c/a and b/a. A small value of c/a suggests two possibilities: if b/a is large, it indicates an oblate distribution resembling a thin plane; if b/a is similarly small ($c/a\approx b/a$), it suggests a narrow, prolate distribution resembling a filament-like shape (Metz et al., 2007).
Figure 2: Comparison of eccentricity $(e)$ distributions across different $\theta$. The figure presents the eccentricity CDF for three different $\theta_{max}$: 40°, 60°, and 80°, from left to right, respectively, shown in gray. The solid black line in each panel represents the observed data for Milky Way satellites from (Li et al., 2021), while the overlaid gray CDFs correspond to simulated results for $N_{\text{sat}}$ test satellites.
@@ 2.3 Adopted potential
Understanding the gravitational potential and mass distribution of the Milky Way is crucial to numerically integrating the orbits of test satellites. Despite extensive research over the past century, the exact distribution of the Milky Way’s mass remains a topic of debate. Current estimates of the total mass of the Milky Way range between $0.5-3\times 10^{12}M_{\odot}$ (Wang et al., 2020), varying based on the methodology and the underlying assumptions (Bland-Hawthorn & Gerhard, 2016). Therefore, significant uncertainty surrounds the mass of the Milky Way, necessitating consideration of these ambiguities in mass distribution.
For this study, we used one fiducial potential for forward integration and three distinct gravitational potentials, as detailed in Table 1, to investigate their effects on the evolution and inferred stability of the plane of satellite galaxies in backward integration. This allowed us to evaluate the effect of adopting an incorrect Milky Way mass on the inferred satellite plane’s evolution. To numerically integrate orbits, we relied on Galpy (Bovy, 2015) to investigate the orbital properties of the test satellites. Within the galpy.potential module, the MWPotential2014 model represents the Milky Way’s gravitational potential and is obtained by fitting to a wide range of observational data. This simplified static potential model comprises the following components:
@@
with a power-law index of $\alpha=1.8$ and a cut-off radius of $r_{c}=1.9$ kpc. The disk component is modeled using the axisymmetric Miyamoto–Nagai disk (Miyamoto & Nagai, 1975):
@@
The dark matter halo is characterized by the NFW halo (Navarro, 1996):
@@
The MWPotential2014 model is scaled such that the circular velocity at a distance of 8 kpc from the galactic center within the disc ($z=0$ kpc) is 220 km/s, with a virial mass of approximately $8\times 10^{11}M_{\odot}$. In this study, we refer to this potential as the fiducial potential, $MW_{\text{fiducial}}$. For the two remaining Milky Way potentials, we adjusted $MW_{\text{fiducial}}$ by rescaling the halo mass to achieve the desired mass distribution. Specifically, the $MW_{\text{low}}$ model was configured with a halo mass 40% lower than that of $MW_{\text{fiducial}}$, while $MW_{\text{high}}$ had a halo mass 40% higher than that of $MW_{\text{fiducial}}$, thereby providing deeper insights into the influence of mismatching the Milky Way’s mass on the inferred behavior and stability of satellite galaxy planes.
Table 2: Parameters for forward integration models
@@
111 The table lists the model numbers (FI), the number of satellites ($N_{\text{sat}}$), the maximum angle $\theta_{max}$ in degrees, the gravitational potential model ($\textit{MW}_{fiducial}$), and the number of realizations ($N_{\text{realization}}$) for each forward-integration scenario.
@@ 2.4 Forward Integration
We conducted forward integration of the satellite system over a 5 Gyr time interval for $N$ = 20 random realizations of the initial setup. We selected three distinct maximum tangential velocity angles, $\theta_{\text{max}}$ = 40°, 60°, and 80°, for forward integration, as shown in Table 2. Notably, larger angles allow for more eccentric orbits.
The resulting cumulative distribution functions (CDFs) of eccentricity for the three different values of $\theta$ are shown in Fig. 2. We also overlaid, as a solid black curve, the eccentricity data for observed Milky Way satellites from (Li et al., 2021) onto the plot, which contains data for 46 satellite galaxies from Fritz et al. (2018), McConnachie & Venn (2020) and Simon (2019). However, for our comparison, we selected 35 satellite galaxies, as 11 had eccentricities $e\gg 1$. This addition allowed us to compare the simulated eccentricities with those of observed Milky Way satellite galaxies. From the comparison, it is apparent that $\theta_{\text{max}}$ = 80°aligns closely with the observed eccentricities of satellite galaxies. Each forward integration was conducted using the $MW_{\text{fiducial}}$ potential model. Figure 2 also shows that the CDF range for Li et al. (2021) spans approximately 0.2 to 1.0. For the simulated eccentricities of our initial test runs, the range was between 0.0 and 1.0, because circular orbits were also included. To align the simulated eccentricities with the observed data, we excluded values ranging between $\theta$ = -20°and 20°to ensure that perfectly circular orbits were not generated.
@@ 2.5 Uncertainties
Figure 3: Effect of proper motion uncertainties on the plane of satellite galaxies. Each panel shows $N_{\text{realization}}$ forward integrations as black curves with the mean indicated by a dot-dashed line, and $M_{\text{realization}}$ backward integrations as green curves with the mean indicated by a dashed line.
To precisely determine the motion of a satellite galaxy, its position, distance, and 3D motion are crucial. Spectroscopy is used to track the motion of a galaxy via the Doppler effect. However, it is essential to note that radial velocity measures only one dimension. To determine how a galaxy moves tangentially across the sky, proper motions are measured. The Hubble Space Telescope (HST) has measured the proper motion of most of the 11 brightest classical satellite galaxies of the Milky Way (Redd, 2018). Since then, largely due to the advent of Gaia, the availability of proper motion measurements has improved significantly (Gaia Collaboration et al., 2016b, a). However, highly accurate proper motion measurements remain limited and are available only for the more massive, luminous, or nearby satellite galaxies.
According to Gaia Data Release 2 (DR2), the random proper motion errors for Milky Way satellites range from approximately 0.006 to 0.641 mas yr-1 (Fritz et al., 2018), and the corresponding systematic errors are estimated to be approximately 0.035 to 0.066 mas yr-1 (Gaia Collaboration et al., 2018a, b). In Gaia Early Data Release 3 (EDR3), the uncertainties in proper motions, derived from both statistical and systematic errors, have approximately halved(Lindegren et al., 2021; McConnachie & Venn, 2020; Li et al., 2021; Battaglia et al., 2022).
In this study, we investigated four different ranges of proper motion uncertainties: $\epsilon_{\mu}=0.00$ mas yr-1, 0.04 mas yr-1, 0.08 mas yr-1, and 0.12 mas yr-1. We specifically opted for 0.00 mas yr-1 proper motion uncertainties to represent cases where proper motion was measured with the highest accuracy. This condition allowed us to examine the evolution of the plane of satellite galaxies under various parameters, independent of proper motion errors. We randomly sampled these proper motion uncertainties from Gaussian distributions with width $\epsilon_{\mu}$.
For distance uncertainties, we chose two specific values: $\epsilon_{\mathrm{dist}}=0$% and 5%. Distances are crucial factors in understanding the spatial distribution and dynamics of satellite galaxies around the Milky Way, as well as their effects on the inferred stability of the plane of satellite galaxies. Again, distance errors were sampled from a Gaussian with width $\epsilon_{\mathrm{dist}}$.
@@ 2.6 Backward integration
After completing the forward integration, we recorded the orbital parameters, such as right ascension, declination, proper motion in right ascension and declination, distance, and line-of-sight velocity, for each test satellite system at the final snapshot of all simulation realizations. We then mocked-observed these systems by adding uncertainties in proper motions and distance such that
@@
$\text{dist}_{\text{new}}$ $=\text{dist}+\epsilon_{\text{dist}}.$ (12)
Equations (10) – (12) provide the updated proper motions and distance parameters for each test satellite, incorporating their respective uncertainties. The values of $\epsilon_{\mu}$ were drawn from a Gaussian distribution. In contrast, for $\epsilon_{\text{dist}}$, a percentage error was drawn from a Gaussian distribution, so that more distant satellites had larger absolute distance uncertainties. Table 4 details the ranges of uncertainties in these parameters. After performing mock observations, we integrated the system backward for 5 Gyr. We repeated this process 30 times for each of the 20 forward integrations, using different randomly generated errors each time, resulting in a total of 600 backward integrations to return the test satellite system to its initial state. We conducted this procedure using all three Milky Way potential models. Each row in Table 4 represents a combination of various parameters that could influence the inferred stability of the satellite plane. Overall, with four proper motion uncertainties, two distance uncertainties, and three potential models, there were 24 model combinations for each $\theta_{max}$. Each model underwent 600 backward integrations, totaling 43,200 integrations in total.
@@ 3 Results
In this section we present the main results of the simulation  [^2] GitHub repository of the code used to generate the results: Plane of Satellite Galaxies. For brevity, and because it most closely resembles the eccentricities of the observed Milky Way satellite galaxies, we show only results for $\theta_{max}=80$. The results for the other $\theta_{\text{max}}$ are qualitatively similar and are provided in Appendix A.
3.1 Effect of proper motion uncertainties
@@
Figure 3 presents results that focus on introducing proper artificial motion and distance uncertainties to the test satellite system before backward integration. These results are divided into four panels, each corresponding to a distinct level of proper motion uncertainty.
@@
Each panel shows an array of curves, in gray and green, indicating the $N_{\text{realization}}$ and $M_{\text{realization}}$ numbers for forward and backward integration, respectively. The average value at each time step is shown as a dashed-dotted curve, with arrows pointing to the right and left. In this case, we adopt $MW_{\text{fiducial}}$ for both forward and backward integrations, without introducing any distance uncertainties. In the upper left panel, when $\epsilon_{\mu}$ = 0.00 mas yr-1, the mean minor-to-major ratio, $c/a$, obtained from backward integration, aligns perfectly well with the mean minor-to-major ratio obtained from forward integration, as each forward integration curve is overplotted by backward integration. This demonstrates a high degree of predictability and confirms that the orbital plane remains stable and unchanged in scenarios where the proper motion is measured with perfect accuracy. It also implies that, at least in the theoretical framework of the simulation, the inherent dynamics of the satellite system are sufficiently deterministic to allow precise backtracking of its orbital path, provided that the initial conditions are precisely known. These results serve as a control scenario against which the other panels, which include proper motion errors, can be compared. Model BI-01 from Table 4 shows similar behavior: even after 3 Gyr of backward integration, the difference between forward and backward, $\Delta_{c/a}=0.0$, and the fractional change, $f_{c/a}$, is 1.00, meaning that both forward and backward integration yield the same output when proper motion uncertainties are zero.
@@
In the upper right panel, we have applied Gaia-level systematic uncertainties $\epsilon_{\mu}$ = 0.04 mas yr-1. Here, a slight uncertainty in proper motion leads to a visible divergence between the mean axial ratios of forward and backward integrations. The mean axial ratio from the backward integration does not align as closely with the forward integration as in the first case, when there were no uncertainties. This suggests a decrease in the predictability and inferred stability of the orbital plane with the introduction of proper motion errors. However, the curve appears to retain its shape to some degree. The effect of uncertainties is also evident from Model BI-02 in Table 4, which shows that with the introduction of $\epsilon_{\mu}=0.04$ mas yr-1, $\Delta_{c/a}$ shifts from 0.00 to 0.073, and $f_{c/a}$ becomes 1.69. Thus, the inferred past plane width is typically 70% higher than the true value.
@@
Increasing the proper motion uncertainties to 0.08 mas yr-1 further increases the divergence between the means of backward and forward integration, as shown in the lower left panel. This indicates a further decline in the predictability and stability of the orbital plane as proper motion uncertainty increases. Moreover, Table 4 also shows that in Model BI-03, $\Delta_{c/a}$ shifts yet again, from 0.07 to 0.17, and $f_{c/a}$ rises to 2.64, meaning that the typical inferred plane width is almost three times larger than the true one.
@@
Under extreme proper motion uncertainties, 0.12 mas yr-1, the divergence between the mean forward and mean backward integration, shown by dashed curves, is more pronounced. The mean axial ratio from backward integration shows a significant divergence from the forward integration mean, implying a severe reduction in the orbital plane’s predictability and stability with the highest proper motion uncertainty examined.
@@
Moreover, the original shape of the curve is no longer discernible and deviates significantly more from the forward integration curve compared to the other uncertainties examined. This suggests that the actual past evolution of the system cannot be reliably traced under such high uncertainties. This shows that the inferred dynamics of the system are highly sensitive to proper motion uncertainties, making it challenging to draw reliable conclusions about the satellites’ orbital evolution under these conditions. This case is represented by Model BI-04 in Table 4, which shows that $\Delta_{c/a}$ = 0.24 and $f_{c/a}$ = 3.32.